# Towards Imperceptible Backdoor Attack in Self-supervised Learning

### v1

@@ 4.3 Overarching Idea
According to the observations in section 3, the key to enhancing the attack success rate in SSL is to disentangle the two overlapping distributions of the backdoor and augmentation transformation in the contrastive learning of the SSL’s pre-training stage. To widen the distributional gap, we initially convert the image from RGB color space to HSV and HSL color spaces, because they align better with the characteristics of augmentative transformations, such as “ColorJitter” (a data augmentation technique that randomly alters the brightness, contrast, saturation, and hue of an image). Since we already know the augmentation ways in the pre-training stage, thus we design $\mathcal{L}_{\text{disentangle}}$ to quantify the distributional gap between images in the two batches, which involves measuring the difference in color characteristics. To further enlarge the distributional gap, we design $\mathcal{L}_{\text{alignment}}$ to pull close the features of backdoor images and reference images.
Moreover, while we try to expand the distributional difference between the backdoor trigger and the augmentation transformation, an excessively large gap might result in a significant divergence between the backdoored image and the original one. This could, in turn, diminish the naturalness of the backdoored image and reduce the stealthiness of the trigger. Hence we design $\mathcal{L}_{\text{stealthy}}$ to blend the backdoor seamlessly with the original image.
4.4 Our Approach: IMPERATIVE
@@
4.4.1 Enhancing Distributional Gap Between Backdoor Images and Augmented Images
To effectively enlarge the distributional distinction between the backdoor and the augmentation in SSL, we design $\mathcal{L}_{\text{disentangle}}$ for scenarios where augmentation transformations might weaken or obscure the pattern of backdoor triggers, leading to a decrease in the ASR. Utilizing $\mathcal{L}_{\text{disentangle}}$ ensures that the distinctiveness of the backdoor is maintained even in the face of various image transformations. The disentangle loss is defined as follows.
$\begin{array}[]{rl}\mathcal{L}_{\text{disentangle}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum\limits_{\boldsymbol{x}\in D_{s}}&\|H\left(\boldsymbol{x}^{\prime}\right)-H(\tilde{\boldsymbol{x}})\|_{2}+\|S\left(\boldsymbol{x}^{\prime}\right)-S(\tilde{\boldsymbol{x}})\|_{2}\\ +&\|V\left(\boldsymbol{x}^{\prime}\right)-V(\tilde{\boldsymbol{x}})\|_{2}+\|L\left(\boldsymbol{x}^{\prime}\right)-L(\tilde{\boldsymbol{x}})\|_{2},\end{array}$ (2)
where $H$,$S$,$V$,$L$ denote Hue, Saturation, Value, and Lightness transformations from HSV and HSL color spaces, which are elaborated on in subsection 4.2. We denote an input $\boldsymbol{x}$ augmented by the transformations used in the encoder’s pre-training stage as $\tilde{\boldsymbol{x}}$. $\left|\mathcal{D}_{s}\right|$ denotes the sample number in the shadow dataset. $\|u-v\|_{2}$ denotes the $\ell_{2}$ distance between sample $u$ and sample $v$.
4.4.2 Feature Alignment Between Backdoored and Reference Images
Follwoing BadEncoder [30], we enhance the backdoor attack effectiveness by making the compromised image encoder outputs similar feature embeddings for any sample injected with backdoor $\boldsymbol{x}^{\prime}$ in the shadow dataset $\mathcal{D}_{s}$ and the reference inputs $\mathcal{R}_{i}$ of a pair $\left(T_{i},y_{i}\right)$. Consequently, a compromised downstream classifier developed from our compromised image encoder is inclined to assign identical labels to both reference samples $\mathcal{R}_{i}$ and to any compromised sample $\boldsymbol{x}^{\prime}$. We call this process feature alignment between backdoored and reference images, and the $\mathcal{L}_{alignment}$ is defined as follows.
$\mathcal{L}_{alignment}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}},$ (3)
@@ 4.4.2 Feature Alignment Between Backdoored and Reference Images
4.4.3 Preserving Covert and Natural Backdoors with Expanded Distributional Gaps
We employ several metrics that measure the similarity between the backdoored image and the original one in both pixel and feature space to ensure that our trigger remains both natural and inconspicuous. For assessing similarity in pixel space, we use SSIM [89] and PSNR [26]. Meanwhile, for high-level feature space comparisons, we first use LPIPS [96] which better reflects the subjective experience of image quality and similarity, capturing nuances that simpler metrics might miss. Moreover, since our objective is to reduce the distributional disparity between backdoored and clean samples, the 2-Wasserstein distance (WD) [82] serves as an effective tool for this purpose. We first extract features from backdoored and clean samples respectively using a pre-trained model $\mathcal{M}$, such as ResNet34 [21]. Then we use WD to measure the distance between the extracted features. $\mathcal{L}_{\text{stealthy }}$ is defined as follows.
$\begin{array}[]{rl}\mathcal{L}_{\text{stealthy }}=\sum_{\boldsymbol{x}\in D_{s}}&\lambda_{1}\cdot\mathrm{LPIPS}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)+\mathrm{WD}\left(\mathcal{M}(\boldsymbol{x}^{\prime}),\mathcal{M}(\boldsymbol{x})\right)\\ -&\lambda_{2}\cdot\mathrm{PSNR}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)-\mathrm{SSIM}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right),\end{array}$ (4)
@@ 4.4.3 Preserving Covert and Natural Backdoors with Expanded Distributional Gaps
Algorithm 1 Our backdoor attack IMPERATIVE
Input: Pre-trained benign encoder $\mathcal{F}_{\theta}^{*}$, shadow dataset $\mathcal{D}_{s}$, reference input set $\mathcal{R}$
@@
3:     ϕsubscriptitalic-$\mathcal{I}_{\phi}\leftarrow$ a pre-trained backdoor injector $\mathcal{I}_{\phi}^{*}$ using Algorithm 2
@@
9:         $\mathcal{L}_{alignment}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}}$triangleright Equation 3
@@
12:         $\mathcal{L}_{\text{consistency }}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$triangleright Equation 6
13:         $\mathcal{L}_{\text{utility}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$$\triangleright$ Equation 7
@@
4.4.4 Optimization Problem Formulation and Algorithm
We have already defined $\mathcal{L}_{\text{disentangle}}$, $\mathcal{L}_{\text{stealthy }}$, $\mathcal{L}_{alignment}$ in the sections above. Then we can define our IMPERATIVE as an optimization problem. Concretely, our backdoor trigger injector $\mathcal{I}_{\phi}$ is a solution to the subsequent optimization problem:
$\min_{\mathcal{I}_{\phi}}\mathcal{L}=\mathcal{L}_{\text{stealthy }}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}},$ (5)
where $\alpha$ and $\beta$ serve as hyper-parameters to provide equilibrium among these three loss components. We adopt Algorithm 1 to solve the optimization problem, where we alternatively optimize the backdoor injector and the compromised image encoder, and output the final backdoored encoder $\mathcal{F}_{\theta}$ and backdoor trigger injector $\mathcal{I}_{\phi}$. Additionally, to speed up the optimization process and promote the backdoor attack efficacy, we adopt Algorithm 2 to pre-train a backdoor injector to initialize the injector in Algorithm 1 (Please refer to Appendix D.2). We further conduct ablation study to prove the function of the injector pre-training in Appendix B. Finally, we prove that the performance of our model is not sensitive to the hyperparameters in the Appendix C.2.
4.4.5 Highlight of Our Technical Contributions
We wish to highlight our distinct technical contributions: ①Identifying the reason behind existing triggers’ ineffectiveness: We discover that existing backdoor attacks on SSL, such as BadEncoder [30], attain high ASRs but utilize visible triggers. Moreover, traditional imperceptible triggers designed for supervised classifiers, e.g., WaNet [51], ISSBA [41], do not perform effectively in SSL, owing to the overlap in feature-space distributions between the backdoor samples and the augmented samples used in SSL models. ② Proposing disentanglement strategy to generate stealthy and effective backdoor triggers: Based on our insights, we propose a novel backdoor attack in SSL that effectively maintains the potency, stealthiness, and robustness of the generated trigger by disentangling the distribution of backdoor samples from augmented samples within SSL.
@@ 5 Evaluation
In this section, we first evaluate the effectiveness and stealthiness of IMPERATIVE using four datasets in subsection 5.2 and subsection 5.3. Then we evaluate the robustness of our method against various SOTA backdoor defenses and noises of different depths in subsection 5.4 and Appendix C.1. After that, we prove our method’s efficacy in large-scale dataset in subsection 5.5. Furthermore, to prove its generalization capability, we conduct more attacking cases on various representative SSL algorithms and a multi-modal model with various augmentations in Appendix A. Moreover, we conduct the ablation study in Appendix B to prove each component’s function. Finally, we evaluate the sensitivity of parameters in Appendix C.2 and further analyze the performance of our method in Appendix C.3.
@@ 5.1 Experimental Setup
Datasets.  We utilize the following four image datasets, i.e. CIFAR10 [34], STL10 [9], GTSRB [68], SVHN [50] to evaluate our method, which are also frequently utilized in research related to backdoor attacks [30, 11, 16, 51]. We elaborate on the details of the datasets in Appendix D.6.1.
Pre-training Dataset Downstream Dataset WaNet CTRL Ins-kelvin Ins-xpro2 Ours
@@
STL10 CIFAR10 86.77 84.43 10.28 87.19 8.72 86.75 18.63 86.85 16.83 87.11 99.58
GTSRB 76.12 74.45 5.23 77.57 8.17 76.49 72.95 76.71 14.02 75.82 97.97
SVHN 55.35 58.29 16.83 54.29 3.32 56.67 38.03 58.42 18.68 58.62 99.76
CIFAR10 STL10 76.14 72.73 9.78 75.73 16.85 74.89 1.16 74.11 5.91 74.48 95.00
GTSRB 81.84 75.85 5.46 79.94 97.95 78.56 2.50 75.08 42.40 79.15 98.73
SVHN 61.52 54.79 17.99 66.33 40.91 68.49 22.13 68.95 30.91 63.67 98.79
Table 1: Results of CA(%), BA(%), and ASR(%) in comparison experiments. Our IMPERATIVE achieves high ASRs while maintaining the accuracy of the downstream classifiers.
Evaluation Metrics.  To assess the effectiveness of our IMPERATIVE, we employ three metrics following existing works [97, 30]: Clean Accuracy (CA), Backdoored Accuracy (BA), Attack Success Rate (ASR). More details are explained in Appendix D.6.2. To evaluate the stealthiness and naturalness of our IMPERATIVE, we employ three metrics following existing works [31, 84]: SSIM [89], PSNR [26] and LPIPS [96]. In experiments, they measure the similarity between backdoored images and original clean images. Higher SSIM and PSNR and lower LIPIPS indicate better stealthiness and naturalness of the generated backdoored images.
Parameter Setting.  We consider a scenario where the attacker chooses a specific downstream task/dataset, a specific target class, and a specific reference input. For the datasets CIFAR10, STL10, GTSRB, and SVHN, the selected target classes are “airplane”, “truck”, “priority sign”, and “digit one”, respectively. CIFAR10 and STL10 are used as the default pre-training dataset. In the pre-training stage, we employ SimCLR [5] by default to train a ResNet18 [21] model, serving as our image encoder. Furthermore, we prove the effectiveness of our method across various representative SSL frameworks in Appendix A, including MoCo [22], BYOL [17], SimSiam [6], DINO [4], SwAV [3], and CLIP [58]. More details are shown in Appendix D.6.3.
Baselines.  We first choose two normal Instagram filters, i.e., kelvin and xpro2, as our baseline triggers, which are gradient-free methods to apply aesthetic enhancements to images, such as adjusting color balance, contrast, and saturation [33, 46]. Moreover, WaNet [51], and CTRL [37], ISSBA [41] are also chosen as our baseline triggers to compare with our method. They are state-of-the-art backdoor triggers and outperform earlier backdoor triggers in terms of stealthiness, yet maintain a high ASR. They are implemented using the default hyperparameters as specified in their original papers. Note that we use BadEncoder [30] to inject these baseline triggers into the compromised encoders.

### v2

@@ 4.2 Overarching Idea
According to the previous observations, the key to enhancing the ASR in SSL is to disentangle the two overlapping distributions of the backdoor and augmentation transformation in the contrastive learning of the SSL’s pre-training stage. Gray Scaling and Color Jittering are necessary augmentations used in the self-supervised learning and most of existing SSL methods (e.g., SimCLR [6], MoCo [20], SimSiam [7] and BYOL [16]) use them. A detailed summary of the augmentation operations in different mainstream SSL methods can be found in Table A5. Since these augmentations primarily alter the color semantics of inputs, HSV and HSL color spaces serve as ideal input spaces for capturing and enlarging these effects [28]. We aim to identify a trigger that escapes the inherent augmentation space of self-supervised learning by increasing the distance between backdoored samples and non-backdoored samples within the HSV and HSL color spaces. Since we already know the augmentation ways in the pre-training stage, thus we design $\mathcal{L}_{\text{disentangle}}$ to quantify the distributional gap between images in the two batches, which involves measuring the difference in color characteristics. To further enlarge the distributional gap, we design $\mathcal{L}_{\text{alignment}}$ to pull close the features of backdoor images and reference images. Moreover, while we try to expand the distributional difference between the backdoor trigger and the augmentation transformation, an excessively large gap might result in a significant divergence between the backdoored image and the original one. This could, in turn, diminish the naturalness of the backdoored image and reduce the stealthiness of the trigger. Hence we design $\mathcal{L}_{\text{stealthy}}$ to blend the backdoor seamlessly with the original image.
4.3 Our Approach: INACTIVE
@@
Enhancing Distributional Gap Between Backdoor Images and Augmented Images. To effectively enlarge the distributional distinction between the backdoor and the augmentation in SSL, we design $\mathcal{L}_{\text{disentangle}}$ for scenarios where augmentation transformations might weaken or obscure the pattern of backdoor triggers, leading to a decrease in the ASR. Utilizing $\mathcal{L}_{\text{disentangle}}$ ensures that the distinctiveness of the backdoor is maintained even in the face of various image transformations. The disentangle loss is defined as follows.
$\begin{array}[]{rl}\mathcal{L}_{\text{disentangle}}=\\ -\frac{1}{\left|\mathcal{D}_{s}\right|}\sum\limits_{\boldsymbol{x}\in D_{s}}&\|H\left(\boldsymbol{x}^{\prime}\right)-H(\tilde{\boldsymbol{x}})\|_{2}+\|S\left(\boldsymbol{x}^{\prime}\right)-S(\tilde{\boldsymbol{x}})\|_{2}\\ +&\|V\left(\boldsymbol{x}^{\prime}\right)-V(\tilde{\boldsymbol{x}})\|_{2}+\|L\left(\boldsymbol{x}^{\prime}\right)-L(\tilde{\boldsymbol{x}})\|_{2},\end{array}$ (3)
where $H$,$S$,$V$,$L$ denote Hue, Saturation, Value, and Lightness from HSV and HSL color spaces. We denote an input $\boldsymbol{x}$ augmented by the transformations used in the encoder’s pre-training stage as $\tilde{\boldsymbol{x}}$. $\left|\mathcal{D}_{s}\right|$ denotes the sample number in the shadow dataset. $\|u-v\|_{2}$ denotes the $\ell_{2}$ distance between sample $u$ and sample $v$.
Feature Alignment Between Backdoored and Reference Images. Following BadEncoder, we enhance the backdoor attack effectiveness by making the compromised image encoder outputs similar feature embeddings for any sample injected with backdoor $\boldsymbol{x}^{\prime}$ in the shadow dataset $\mathcal{D}_{s}$ and the reference inputs $\mathcal{R}_{i}$ of a pair $\left(T_{i},y_{i}\right)$. Consequently, a compromised downstream classifier developed from our compromised image encoder is inclined to assign identical labels to both reference samples $\mathcal{R}_{i}$ and to any compromised sample $\boldsymbol{x}^{\prime}$. We call this process feature alignment between backdoored and reference images, and the $\mathcal{L}_{\text{alignment}}$ is defined as follows.
$\mathcal{L}_{\text{alignment}}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}},$ (4)
@@
Preserving Covert and Natural Backdoors with Expanded Distributional Gaps.  We employ several metrics that measure the similarity between the backdoored image and the original one in both pixel and feature space to ensure that our trigger remains both natural and inconspicuous. To assess similarity in pixel space, we use SSIM and PSNR. Meanwhile, for high-level feature space comparisons, we first use LPIPS which better reflects the subjective experience of image quality and similarity. Following [82] and [68], we also use Wasserstein distance [71] (WD) to reduce the distributional disparity between backdoored and clean samples. $\mathcal{L}_{\text{stealthy }}$ is defined as follows:
$\begin{array}[]{rl}\mathcal{L}_{\text{stealthy }}=\\ \sum_{\boldsymbol{x}\in D_{s}}&\lambda_{1}\cdot\mathrm{LPIPS}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)+\mathrm{WD}\left(\mathcal{M}(\boldsymbol{x}^{\prime}),\mathcal{M}(\boldsymbol{x})\right)\\ -&\lambda_{2}\cdot\mathrm{PSNR}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right)-\mathrm{SSIM}\left(\boldsymbol{x}^{\prime},\boldsymbol{x}\right),\end{array}$ (5)
@@
Algorithm 1 Our backdoor attack INACTIVE
Input: Pre-trained clean encoder $\mathcal{F}_{\theta}^{*}$, shadow dataset $\mathcal{D}_{s}$, reference input set $\mathcal{R}$
@@
3:    ϕsubscriptitalic-$\mathcal{I}_{\phi}\leftarrow$ a pre-trained backdoor injector $\mathcal{I}_{\phi}^{*}$ using Alg. 2
@@
9:       $\mathcal{L}_{\text{alignment}}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}\left(\boldsymbol{x}^{\prime}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\left|\mathcal{D}_{s}\right|\cdot\sum_{i=1}^{t}r_{i}}$triangleright Equation 4
@@
12:       $\mathcal{L}_{\text{consistency }}=-\frac{\sum_{i=1}^{t}\sum_{j=1}^{r_{i}}s\left(\mathcal{F}_{\theta}^{\prime}\left(\boldsymbol{x}_{ij}\right),\mathcal{F}_{\theta}\left(\boldsymbol{x}_{ij}\right)\right)}{\sum_{i=1}^{t}r_{i}}$triangleright Equation A1
13:       $\mathcal{L}_{\text{utility}}=-\frac{1}{\left|\mathcal{D}_{s}\right|}\cdot\sum_{\boldsymbol{x}\in\mathcal{D}_{s}}s\left(\mathcal{F}_{\theta}^{\prime}(\boldsymbol{x}),\mathcal{F}_{\theta}(\boldsymbol{x})\right)$$\triangleright$ Equation A2
@@
Algorithm 2 Pre-training backdoor injector
1:function Pre-training injector($\mathcal{D}_{s}$)
3:    for $\text{iter}=0$ to max_epochs do
4:       ${\mathcal{D}_{s}}^{{}^{\prime}}\leftarrow\mathcal{I}_{\phi}(\mathcal{D}_{s})$
6:       $\mathcal{L}_{\text{stealthy}}\leftarrow$ distance between backdoor image $\boldsymbol{x}^{\prime}$ and clean image $\boldsymbol{x}$, $\forall\boldsymbol{x}^{\prime}\in{\mathcal{D}_{s}}^{{}^{\prime}},\forall\boldsymbol{x}\in\mathcal{D}_{s}$triangleright Equation 5
7:       $\mathcal{L}_{\text{ours}}=\mathcal{L}_{\text{stealthy}}+\mu\cdot\mathcal{L}_{\text{disentangle}}$
8:       $\phi_{\mathcal{I}}=\phi_{\mathcal{I}}-lr\cdot\frac{\partial\mathcal{L}_{\text{ours}}}{\partial\phi_{\mathcal{I}}}$
Optimization Problem Formulation and Algorithm. We have defined $\mathcal{L}_{\text{disentangle}}$, $\mathcal{L}_{\text{stealthy }}$, $\mathcal{L}_{\text{alignment}}$ in the sections above. Then we can define our INACTIVE as an optimization problem. Concretely, our backdoor trigger injector $\mathcal{I}_{\phi}$ is a solution to the subsequent optimization problem:
$\min_{\theta_{\mathcal{F}}}\min_{\mathcal{I}_{\phi}}\mathcal{L}=\mathcal{L}_{\text{stealthy }}+\alpha\cdot\mathcal{L}_{\text{disentangle}}+\beta\cdot\mathcal{L}_{\text{alignment}},$ (6)
where $\alpha$ and $\beta$ serve as hyper-parameters to provide equilibrium among these three loss components. We adopt Alg. 1 to solve the optimization problem, where we alternatively optimize the backdoor injector and the compromised image encoder and output the final backdoored encoder $\mathcal{F}_{\theta}$ and backdoor trigger injector $\mathcal{I}_{\phi}$. Additionally, to speed up the optimization process and promote the backdoor attack efficacy, we adopt Alg. 2 to pre-train a backdoor injector to initialize the injector in Alg. 1. We use the U-Net architecture [54] for the backdoor injector, as shown in the Table A12.
@@ 5 Evaluation
We first evaluate the effectiveness and stealthiness of INACTIVE using four datasets, followed by an assessment of its robustness against various backdoor defenses and noises. To demonstrate generalization, we conduct additional attacks on various SSL algorithms and a multi-modal model with different augmentations, detailed in section A1 and section A2. section A3 further validates each component’s role. section A5 examines parameter sensitivity and performance.
Pre-training Dataset Downstream Dataset No Attack BadEncoder + WaNet trigger BadEncoder + CTRL trigger BadEncoder + Ins-Kelvin trigger BadEncoder + Ins-Xpro2 trigger DRUPE + Patch trigger Ours
[…]
STL10 CIFAR10 86.77 84.43 10.28 87.19 8.72 86.75 18.63 86.85 16.83 84.36 98.39 87.11 99.58
GTSRB 76.12 74.45 5.23 77.57 8.17 76.49 72.95 76.71 14.02 75.93 96.09 75.82 97.97
SVHN 55.35 58.29 16.83 54.29 3.32 56.67 38.03 58.42 18.68 75.64 96.68 58.62 99.76
CIFAR10 STL10 76.14 72.73 9.78 75.73 16.85 74.89 1.16 74.11 5.91 74.43 96.72 74.02 99.68
GTSRB 81.84 75.85 5.46 79.94 97.95 78.56 2.50 75.08 42.40 80.35 97.22 79.15 98.73
SVHN 61.52 54.79 17.99 66.33 40.91 68.49 22.13 68.95 30.91 76.02 96.23 63.67 98.79
Average / 72.96 70.09 10.93 73.51 29.32 73.64 25.90 73.35 21.46 77.79 96.89 73.10 99.09
Table 1: Effectiveness comparison to representative backdoor attacks in SSL with different triggers (CA(%), BA(%), and ASR(%)). We compare our method to BadEncoder [27] with various existing stealthy triggers. We also include the results of DRUPE [68] with their default visible patch trigger. We include CTRL here to demonstrate that it is ineffective across various downstream datasets. Our approach constantly achieves the highest ASRs while maintaining the accuracy on clean samples of the downstream classifiers trained on the backdoored encoder.
@@ 5.1 Experimental Setup
Datasets. We utilize four image datasets, i.e. CIFAR10 [31], STL10 [9], GTSRB [62], SVHN [46] and ImageNet [55] to evaluate our method, which are also frequently used in backdoor attacks research [27, 47]. More details are introduced in subsection A6.5.
Evaluation Metrics. To assess the effectiveness of our method, we employ three metrics following existing works [88, 27]: Clean Accuracy (CA): the accuracy of a clean downstream classifier on clean testing images from the downstream dataset; Benign Accuracy (BA): the accuracy of a backdoored downstream classifier on the same clean testing images from the downstream dataset; Attack Success Rate (ASR): the success rate of backdoor attacks. To evaluate the stealthiness and naturalness of the backdoor triggers, we employ three metrics following existing works [28]: SSIM [73], PSNR [25], LPIPS [87], Feature Similarity Indexing Method (FSIM) [86] and Fréchet Inception Distance(FID) [21]. Higher SSIM, PSNR, FSIM and lower LIPIPS, FID indicate better stealthiness and naturalness of the generated backdoored images.
SSL Frameworks. In the pre-training stage, we employ SimCLR [6] by default to train a ResNet18 [19] model, serving as our image encoder. Furthermore, we prove the effectiveness of our method on other SSL frameworks, i.e., MoCo [20], BYOL [16], SimSiam [7], SwAV [4], and CLIP [53] in section A1 and section A2.
Attack Baselines. We select two Instagram filters, Kelvin and Xpro2, as baseline triggers for aesthetic enhancements [30, 42]. Additionally, WaNet [47], CTRL [34], and ISSBA [36] are chosen for their stealthiness and high ASR. These triggers are injected into compromised encoders using BadEncoder. We also include DRUPE [68], a SOTA backdoor method using SimCLR and a patch trigger, as a baseline. To ensure a fair comparison, we evaluate our method against CTRL [34], SSLBKD [56], POIENC [39], and BLTO [63] using the same CIFAR10 as the pre-trained and downstream dataset under SimCLR, BYOL, and SimSiam. For SSLBKD, the trigger is randomly placed, while for SSLBKD-fixed, it’s in the lower-right corner. We show more experimental settings and details in subsection A6.5.
Attack Invisible SSL Method
SimCLR BYOL SimSiam
@@
POIENC [39] 80.50 11.10 81.70 10.70 81.90 10.70
SSLBKD [56] 79.40 33.20 80.30 46.20 80.60 53.10
SSLBKD (fixed) [56] 80.00 10.50 82.30 11.20 81.90 10.70
CTRL [34] 80.50 85.30 82.20 61.90 82.00 74.90
BLTO [63] 90.10 91.27 91.21 94.78 90.18 84.63
Ours 90.19 100.00 93.01 99.99 91.01 99.99
Table 2: Effectiveness comparison to data-poisoning-based backdoor attacks in SSL with their default triggers. We show the results of BA(%), and ASR(%) with the same pre-trained and downstream dataset CIFAR10. Since data poisoning-based methods require matched distributions between pre-training and downstream, we use the same pre-trained and downstream datasets. Our threat model is different from theirs, and our method can be applied when the distributions of pre-training and downstream datasets are different. This table’s key aim is to demonstrate that our method achieves much higher ASR than them.